# A 106K Multi-Topic Multilingual Conversational User Dataset with Emoticons

### v1

@@ Abstract.
Instant messaging has become a predominant form of communication, with texts and emoticons enabling users to express emotions and ideas efficiently. Emoticons, in particular, have gained significant traction as a medium for conveying sentiments and information, leading to the growing importance of emoticon retrieval and recommendation systems. However, one of the key challenges in this area has been the absence of datasets that capture both the temporal dynamics and user-specific interactions with emoticons, limiting the progress of personalized user modeling and recommendation approaches. To address this, we introduce the emoticon dataset, a comprehensive resource that includes time-based data along with anonymous user identifiers across different conversations. As the largest publicly accessible emoticon dataset to date, it comprises 22K unique users, 370K emoticons, and 8.3M messages. The data was collected from a widely-used messaging platform across 67 conversations and 720 hours of crawling. Strict privacy and safety checks were applied to ensure the integrity of both text and image data. Spanning across 10 distinct domains, the emoticon dataset provides rich insights into temporal, multilingual, and cross-domain behaviors, which were previously unavailable in other emoticon-based datasets. Our in-depth experiments—both quantitative and qualitative—demonstrate the dataset’s potential in modeling user behavior and personalized recommendation systems, opening up new possibilities for research in personalized retrieval and conversational AI. The dataset is freely accessible  [^1] https://huggingface.co/datasets/metchee/u-sticker.
@@
copyright: noneconference: Preprint; 2025; $c$s: Information systems Information retrievalccs: Information systems Dataset Personalization
@@ 1. Introduction
Figure 1. The colorful and vivid emoticons makes for an expressive and comfortable conversation.
Instant messaging (IM) has become an essential mode of communication, allowing users to interact efficiently. Beyond text, colorful and animated emoticons have emerged as a powerful medium of expression, enabling users to convey semantics and emotions comfortably and accurately (Ge, 2020; Constantin et al., 2019; Kariko and Anasih, 2019; Gygli and Soleymani, 2016; Jiang et al., 2023), as seen in Figure 1. Unlike general image datasets, which emphasize object recognition, emoticons carry rich semantic meaning, expressiveness, and emotional nuance, making their retrieval and recommendation fundamentally different from traditional image retrieval and recommendation tasks.
However, despite emoticons’ widespread use, personalized emoticon retrieval and recommendation remains an understudied area, primarily due to the lack of large-scale user emoticon interaction dataset. As shown in Table 1, while several emoticon dataset exist, they often lack user information, making them unable to identify the same user across different conversations. Additionally, they contain few emoticons per conversation, providing limited insight into user preferences or emoticon usage patterns. Furthermore, some datasets are not open-sourced, limiting reproducibility and progress of personalized emoticon retrieval and recommendation.
Table 1. Comparison with Emoticon Datasets. We bold the largest count among publicly available datasets and underline the second largest. Our dataset is cross-domain, large in scale and contain time and user ID for personalization.
Dataset Pub. Avail. Cross Domain Time Info. # Sticks./user # Emoticons Description
StickerTag (Wang et al., 2024) 13,571 Emoticon-tag pairs
StickerInt (Liang et al., 2024) Unknown 1,025 Dialogues with emoticons
StickerCLIP (Zhao et al., 2023) 820,000 Chinese emoticon-tag pairs
CSMSA (Ge et al., 2022) 16,000 Emoticons-tag pairs
SRS, PESRS (Gao et al., 2021, 2020) 6.82 320,168 Dialogues with emoticons
PerSRV (Chee et al., 2024) Unknown 543,098 Emoticon-query pairs
MCDSCS (Shi and Kong, 2024) 2.03 14,400 Dialogues with emoticons
SER30K (Liu et al., 2022) 30,739 Emoticon-sentiments pairs
MOD (Fei et al., 2021) 1.36 307 Dialogues with emoticons
Ours 16.90 370,222 Dialogues with emoticons
To address these limitations, we introduce a new emoticon dataset, the first large-scale dataset that includes both user information and emoticon-based conversations. Our contributions can be summarized as follows:
  • We present a new emoticon dataset, the largest emoticon dataset to date, containing 22k users, 370.2k emoticons and 8.3M conversation messages of texts and emoticons.
  • Our dataset is a multi-domain dataset that includes rich and diverse information. Covering 10 domains, it captures temporal, multilingual, and cross-domain behaviors that are not present in previous datasets.
  • Extensive quantitative and qualitative experiments demonstrate the dataset’s practical applications on user behavior analysis and modeling, personalized emoticon recommendation. It also holds potential for further research in areas such as personalized retrieval and conversational studies.
@@ 2. Related Work
We outline the current emoticon datasets and their applications. The terms sticker, emoticon, memes, gif mean approximately the same, dynamic or static multimedia that can be shared to express emotions. Herein, we use the word emoticon.
2.1. Emoticon Datasets
Due to the challenges in preparing emoticon conversation-based datasets, their availability remains limited, particularly for publicly accessible datasets. However, recent research on emoticons has gained more attention. Currently, there are only nine related datasets, with only three being publicly available. A summary of these datasets is provided in Table 1.
@@ 2.1.1. Availability
Many of these datasets are closed-source. Specifically, six datasets are not publicly accessible (Wang et al., 2024; Liang et al., 2024; Zhao et al., 2023; Ge et al., 2022; Gao et al., 2021; Chee et al., 2024), meaning they are either closed-source or require additional access that is often unattainable. This lack of open access presents a significant challenge for emoticon-based research.
2.1.2. Multi Topic Information
Furthermore, none of the aforementioned datasets contain cross-domain information, as the conversations or dialogues are not identified with specific topics. This limitation hinders the exploration of cross-domain emoticon usage, which could provide valuable insights.
@@ 2.1.3. User Information
Additionally, publicly available datasets are scarce in terms of user data, with 1.36 (Fei et al., 2021) and 2.03 (Shi and Kong, 2024) historical emoticons per user. This lack of data makes it difficult to perform user analysis and personalization-related tasks. Our dataset bridges this gap, providing the highest emoticon history per user, with an impressive 16.90 emoticons.
@@ 2.1.4. Multilingualism
Most existing datasets focus on one or two languages, limiting the ability of smaller communities to benefit from advancements in emoticon personalization. In contrast, our dataset is multilingual, allowing a broader range of communities to engage with these advancements.
@@
As shown in Table 1 and our analysis, our dataset is the largest publicly available emoticon-user dataset. It features 370,222 emoticons, cross-domain behavior, rich contextual information, and comprehensive user and temporal data.
2.2. Emoticon Retrieval
@@
Most previous research emphasizes the importance of data for emoticon retrieval. SRS (Gao et al., 2020), PESRS (Gao et al., 2021) require corresponding utterances, while Lao et al. (Lao et al., 2022) rely on manually labeled emotions, sentiments, and reply keywords. CKES (Chen et al., 2024) annotates each emoticon with a corresponding emotion. During emoticon creation, Hike Messager (Laddha et al., 2020) tags conversational phrases to emoticons. The reliance on data presents a significant limitation, as emoticons without associated information are excluded from consideration.
@@
Gao et al. (Gao et al., 2020) use a convolutional emoticon encoder and self-attention dialog encoder for emoticon-utterance representations, followed by a deep interaction network and fusion network to capture dependencies and output the final matching score. The method selects the ground truth emoticon from a pool of emoticon candidates and its successor. Zhang et al. (Zhang et al., 2022) perform this on recommendation tasks. CKES (Chen et al., 2024) introduces a causal graph to explicitly identify and mitigate spurious correlations during training. The PBR (Xia et al., [n. d.]) paradigm enhances emotion comprehension through knowledge distillation, contrastive learning, and improved hard negative sampling to generate diverse and discriminative emoticon representations for better response matching in dialogues. PEGS (Zhang et al., 2024), StickerInt (Liang et al., 2024) generate emoticon information using multimodal models and selects emoticon responses, but does not consider personalization. StickerCLIP (Zhao et al., 2023) fine-tunes pretrained image encoders but does not consider personalization. In addition, many methods designed to rank emoticons from top-k candidates face a significant drawback in real-world emoticon retrieval scenarios, as they quickly become impractical when applied to larger datasets. Possibly due to the lack of dataset, most methods do not consider personalization.
2.3. Personalized Emoticon Recommendation
@@
The absence of personalized emoticon recommendation can likely be attributed to the scarcity of emoticon-user datasets. PESRS (Gao et al., 2021) improves upon earlier work (Gao et al., 2020) by incorporating user preferences, but the dataset it relies on is not publicly accessible. This limited availability indirectly hampers the development of personalized emoticon recommendations and may impede further advancements in the field.
3. Our Dataset
@@
Telegram (tel, [n. d.]) is a widely used open-source messaging platform, offering publicly accessible conversations. The interactions within these groups are natural, making them a valuable source for understanding human behavior, particularly in the context of personalization. As such, they present a promising dataset for emoticon-related research (Dargahi Nobari et al., 2017; Naseri and Zamani, 2019; Baumgartner et al., 2020; Blas et al., 2024; Khaund et al., 2020; La Morgia et al., 2025).
@@ 3.1. Construction Criteria
We firstly define the envision for the final user-emoticon dataset. Since emoticons are the focal point of this dataset, we would like a high presence of emoticons. Additionally, since we would like to analyze interactions across different domains, the final dataset should involve diverse topics. We would also like the dataset to be of high-quality; avoidance of spam, advertisement and groups with single person. Moreover, we noticed that previous studies overlooked linguistic variety, hence, we wish our final dataset to have multilingual coverage. Finally, to ensure a robust and generalizable dataset, we strive for sufficient dataset size.
@@
    Emoticon prevalence: Conversations must contain a significant number of emoticons.
@@
Following these criteria, we manually screened hundreds of conversation groups and ultimately selected 71 conversation groups to crawl their content.
@@ 3.2. Dataset Pre-processing
The dataset, crawled using Telethon (Exo, 2025), contains key components from user messages, including the unique conversation group ID, group name, message datetime, user ID, media types, message ID, message text, and emoticon details (ID, set ID, access hash, and mime type). Users may participate in multiple conversation groups.
@@ 3.2.1. Text Processing
After language detection, we manually review the classification results, removing entries with fewer than 20 occurrences and those from administrative bots. The final language distribution is illustrated in Figure  2. In total, we detect 18 languages, with the top ten languages in order being English, Russian, French, Spanish, Chinese, Polish, Italian and German, Turkish, Portugese. As can be seen, emoticonU features 18 languages which is potentially viable for multilingualism emoticon-related task.
Figure 2. Distribution of detected languages in the dataset. The dataset features over 18 languages. The top ten languages are English, Russian, French, Spanish, Chinese, Polish, Italian and German, Turkish, Portugese.
3.2.2. Emoticon Processing
The downloaded emoticons come in three formats; .webp, .webm, and .tgs, we present the extension distribution in Figure 3. We convert all tgs files to the .gif extension using tgsconverter (Benau, 2025), .webp is converted to .png and no conversion is done to .webm. We break up dymamic images into frames and save a frame every second to prepare for later use.
Figure 3. Preliminary emoticons extension distribution. 65.6% of the emoticons are static and 34.4% are dyanimc.
@@ 3.3.1. Unsafe Text Detection
Offensive text are content that could contain racism, vulgarities or other potentially offensive content. Since different languages require specialized tools for detecting offensive content, we focus on the ten most used languages in our dataset: English, Russian, French, Spanish, Chinese, Polish, Italian and German, Turkish, Portugese. For each, we employ well-established moderation models that have been published in peer-reviewed conferences, we describe the models in the following and the threshold values.
@@
We present the offensive language detection results in the Figure  4, where the blue are safe and red are potentially offensive.
Figure 4. Distribution of offensive comments among language groups.
@@ 3.3.2. Unsafe Text Replacement
  • Offensive Language Replacement: Offensive or harmful language detected in text messages is replaced with the placeholder label: detects this as offensive text. This approach preserves conversation structure while flagging offensive text, ensuring safety without losing context.
  • URL/Link Replacement: When URLs or links are identified in messages, they are replaced with the label: emoticonU detects a link. This ensures private or potentially harmful web addresses are not exposed in the dataset, while preserving the flow of conversation and preventing leakage of sensitive data.
@@ 3.4.1. Unsafe Image Detection
Figure 5. Distribution of nsfw vs. sfw amongst the four offensive image detection models on the raw data.
We utilize the frame processed above and static images into the four models, where we obtain the nsfw to sfw ratio. We present the results in Figure 5. nsfw-image-detection is 0.88%, nsfw-classifier is 0.93%, vit-base-violence-detection is 27.27%, vit-base-nsfw-detector is 0.47%. Then, 471,517 (92.06%) emoticons do not contain offensive taggings and and 40,675 (7.94%) images are considered offensive. Finally, there are a remaining 67 conversation groups.
@@ 3.4.2. Unsafe Image Removal and Replacement
For the identified offensive images, we replace the textual content with the placeholder detects this as offensive image. This helps maintain the integrity of the dataset while ensuring that no harmful visuals, such as explicit or violent content, are retained, thus reducing the risk of privacy violations or inappropriate representations. Then, we remove the emoticon entirely from our emoticon database.
@@ 3.5.1. User Identifier
To protect the user’s privacy, we anonymize the user identifiers by hashing their original integer identifiers using the SHA-256 hash function (Appel, 2015). Originally, the user identifiers were integers ranging from five-digital value to a maximum value of ten-digit value. After applying the SHA-256 hashing algorithm (Appel, 2015), the user identifiers are transformed into 256-bit strings, which are irreversible and do not retain any direct correlation with the original numeric values. This ensures that the users’ identities are kept private, while still allowing for secure and effective data processing and analysis. one more

### v2

# U-Sticker: A Large-Scale Multi-Domain User Sticker Dataset for Retrieval and Personalization

@@ Abstract.
Instant messaging with texts and stickers has become a widely adopted communication medium, enabling efficient expression of user semantics and emotions. With the increased use of stickers conveying information and feelings, sticker retrieval and recommendation has emerged as an important area of research. However, a major limitation in existing literature has been the lack of datasets capturing temporal and user-specific sticker interactions, which has hindered further progress in user modeling and sticker personalization. To address this, we introduce User-Sticker, a dataset that includes temporal and user anonymous ID across conversations. It is the largest publicly available sticker dataset to date, containing 22K unique users, 370K stickers, and 8.3M messages. The raw data was collected from a popular messaging platform from 67 conversations over 720 hours of crawling. All text and image data were carefully vetted for safety and privacy checks and modifications. Spanning 10 domains, the U-Sticker dataset captures rich temporal, multilingual, and cross-domain behaviors not previously available in other datasets. Extensive quantitative and qualitative experiments demonstrate U-Sticker’s practical applications in user behavior modeling and personalized recommendation and highlight its potential to further research areas in personalized retrieval and conversational studies. U-Sticker dataset is publicly available [^1] https://huggingface.co/datasets/metchee/u-sticker.
@@
journalyear: 2025conference: the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13-18, 2025; Padua, Italyjournalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730311isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Information retrievalccs: Information systems Dataset Personalization
@@ 1. Introduction
Instant messaging (IM) has become an essential mode of communication, allowing users to interact efficiently. Beyond text, stickers have emerged as a powerful medium of expression, enabling users to convey semantics and emotions comfortably and accurately (Zhao et al., 2023; Fei et al., 2021). Unlike general image datasets, which emphasize object recognition, stickers carry rich semantic meaning, expressiveness, and emotional nuance, making their retrieval and recommendation fundamentally different from traditional image retrieval and recommendation tasks.
However, despite stickers’ widespread use, personalized sticker retrieval and recommendation remains an understudied area, primarily due to the lack of large-scale user sticker interaction dataset. As shown in Table 1, while several sticker dataset exist, they often lack user information, making them unable to identify the same user across different conversations. Additionally, they contain few stickers per conversation, providing limited insight into user preferences or sticker usage patterns. Furthermore, some datasets are not open-sourced, limiting reproducibility and progress of personalized sticker retrieval and recommendation.
Table 1. Comparison with Stickers Datasets. We bold the largest count among publicly available datasets and underline the second largest. Our dataset is cross-domain, large in scale, and contains time and user ID for personalization.
Dataset Pub. Avail. Cross Domain Time Info. # Sticks./user # Stickers # Languages
StickerTag (Wang et al., 2024) 13,571 1
StickerInt (Liang et al., 2024) Unknown 1,025 1
StickerCLIP (Zhao et al., 2023) 820,000 1
CSMSA (Ge et al., 2022) 16,000 1
SRS, PESRS (Gao et al., 2021, 2020) 6.82 320,168 1
PerSRV (Chee et al., 2024) Unknown 543,098 1
MCDSCS (Shi and Kong, 2024) 2.03 14,400 2
SER30K (Liu et al., 2022) 30,739 1
MOD (Fei et al., 2021) 1.36 307 2
Ours 16.90 370,222 18
To address these limitations, we introduce U-Sticker, the first large-scale dataset that includes both user information and sticker-based conversations. Our contributions are summarized as follows:
  • We present U-Sticker, the largest sticker dataset to date, containing 22K users, 370K stickers and 8.3M conversation messages of texts and stickers.
  • U-Sticker is a multi-domain dataset that includes rich and diverse information. Covering 10 domains, it captures temporal, multilingual, and cross-domain behaviors that are not present in previous datasets.
  • Extensive quantitative and qualitative experiments demonstrate U-Sticker’s practical applications on user behavior analysis and modeling, personalized sticker recommendation. It also holds potential for further research in areas such as personalized retrieval and conversational studies.
@@ 2. Related Work
2.1. Sticker Datasets
Due to the challenges in preparing sticker conversation-based datasets, their availability remains limited, particularly for publicly accessible datasets. However, recent research on stickers has gained more attention. Currently, there are only nine related datasets, with only three being publicly available. A summary of these datasets is provided in Table 1.
@@ 2.1.1. Availability
Many of these datasets are closed-source. Specifically, six datasets are not publicly accessible (Wang et al., 2024; Liang et al., 2024; Zhao et al., 2023; Ge et al., 2022; Gao et al., 2021; Chee et al., 2024), meaning they are either closed-source or require additional access that is often unattainable. This lack of open access presents a significant challenge for sticker-based research.
2.1.2. Cross Domain Information
Furthermore, none of the aforementioned datasets contain cross-domain information, as the conversations or dialogues are not identified with specific topics. This limitation hinders the exploration of cross-domain sticker usage, which could provide valuable insights.
@@ 2.1.3. User Information
Additionally, publicly available datasets are scarce in terms of user data, with 1.36 (Fei et al., 2021) and 2.03 (Shi and Kong, 2024) historical stickers per user. This lack of data makes it difficult to perform user analysis and personalization-related tasks. U-Sticker bridges this gap, providing the highest sticker history per user, with an impressive 16.90 stickers.
@@ 2.1.4. Multilingualism
Most existing datasets focus on one or two languages, limiting the ability of smaller communities to benefit from advancements in sticker personalization. In contrast, U-Sticker is multilingual, allowing a broader range of communities to engage with these advancements.
@@
As see in Table 1 and above, U-Sticker is the largest available sticker-user dataset. It features 370,222 stickers, cross-domain behavior, rich contextual information, and comprehensive user and temporal data.
2.2. Sticker Retrieval
@@
Most previous research emphasizes the importance of data for sticker retrieval. SRS (Gao et al., 2020), PESRS (Gao et al., 2021) require corresponding utterances, while Lao et al. (Lao et al., 2022) rely on manually labeled emotions, sentiments, and reply keywords. CKES (Chen et al., 2024)annotates each sticker with a corresponding emotion. During sticker creation, Hike Messager (Laddha et al., 2020) tags conversational phrases to stickers. The reliance on data presents a significant limitation, as stickers without associated information are excluded from consideration.
@@
Gao et al. (Gao et al., 2020) use a convolutional sticker encoder and self-attention dialog encoder for sticker-utterance representations, followed by a deep interaction network and fusion network to capture dependencies and output the final matching score. The method selects the ground truth sticker from a pool of sticker candidates and its successor. Zhang et al. (Zhang et al., 2022) perform this on recommendation tasks. CKES (Chen et al., 2024) introduces a causal graph to explicitly identify and mitigate spurious correlations during training. The PBR (Xia et al., [n. d.]) paradigm enhances emotion comprehension through knowledge distillation, contrastive learning, and improved hard negative sampling to generate diverse and discriminative sticker representations for better response matching in dialogues. PEGS (Zhang et al., 2024), StickerInt (Liang et al., 2024) generate sticker information using multimodal models and selects sticker responses, but does not consider personalization. StickerCLIP (Zhao et al., 2023) fine-tunes pretrained image encoders but does not consider personalization. In addition, many methods designed to rank stickers from top-k candidates face a significant drawback in real-world sticker retrieval scenarios, as they quickly become impractical when applied to larger datasets. Possibly due to the lack of dataset, most methods do not consider personalization.
2.3. Personalized Sticker Recommendation
@@
The absence of personalized sticker recommendation can likely be attributed to the scarcity of sticker-user datasets. PESRS (Gao et al., 2021) improves upon earlier work (Gao et al., 2020) by incorporating user preferences, but the dataset it relies on is not publicly accessible. This limited availability indirectly hampers the development of personalized sticker recommendations and may impede further advancements in the field.
3. U-Sticker Dataset
@@
Telegram is a widely used open-source messaging platform, offering publicly accessible conversations. The interactions within these groups are natural, making them a valuable source for understanding human behavior, particularly in the context of personalization. As such, they present a promising dataset for sticker-related research.
@@ 3.1. Construction Criteria
We firstly define the envision for the final user-sticker dataset. Since stickers are the focal point of this dataset, we would like a high presence of stickers. Additionally, since we would like to analyze interactions across different domains, the final dataset should involve diverse topics. We would also like the dataset to be of high-quality; avoidance of spam, advertisement and groups with single person. Moreover, we noticed that previous studies overlooked linguistic variety, hence, we wish our final dataset to have multilingual coverage. Finally, to ensure a robust and generalizable dataset, we strive for sufficient dataset size.
@@
    Sticker prevalence: Conversations must contain a significant number of stickers.
@@
Following these criteria, we manually screened hundreds of conversation groups and ultimately selected 70 conversation groups to crawl their content.
@@ 3.2. Dataset Pre-processing
We crawl the data using Telethon (Exo, 2025) and process the data in the following sections. Note that users can participate in multiple conversation groups.
@@ 3.2.1. Text Processing
After language detection, we manually review the classification results, removing entries with fewer than 20 occurrences and those from administrative bots. In total, we detect 18 languages, with the top ten languages in order being English, Russian, French, Spanish, Chinese, Polish, Italian and German, Turkish, Portugese. As can be seen, U-Sticker features 18 languages which is potentially viable for multilingualism sticker-related task.
3.2.2. Sticker Processing
The downloaded stickers come in three formats; .webp, .webm, and .tgs, the distribution is 65.6%, 23.0% and 11.4% respectively. We convert all tgs files to the .gif extension using tgsconverter (Benau, 2025), .webp is converted to .png and no conversion is done to .webm. We break up dymamic images into frames and save a frame every second to prepare for later use.
@@ 3.3.1. Unsafe Text Detection
Offensive text are content that could contain racism, vulgarities or other potentially offensive content. Since different languages require specialized tools for detecting offensive content, we focus on the ten most used languages in our dataset: English, Russian, French, Spanish, Chinese, Polish, Italian and German, Turkish, Portugese. For each, we employ well-established moderation models that have been published in peer-reviewed conferences, we describe the models in the following and the threshold values. The threshold values were manually selected through stepwise examination of the model outputs, aiming to balance the inclusion of high-quality text with the exclusion of plausibly offensive content.
@@
We present the offensive language detection results in Figure  1.
Figure 1. Offensive comment distribution across text classification models. 0.43% of comments are flagged as offensive.
@@ 3.3.2. Unsafe Text Replacement
  • Offensive Language Replacement: Offensive or harmful language detected in text messages is replaced with the placeholder label: U-Sticker detects this as offensive text. This approach preserves conversation structure while flagging offensive text, ensuring safety without losing context.
  • URL/Link Replacement: When URLs or links are identified in messages, they are replaced with the label: U-Sticker detects a link. This ensures private or potentially harmful web addresses are not exposed in the dataset, while preserving the flow of conversation and preventing leakage of sensitive data.
@@ 3.4.1. Unsafe Image Detection
Figure 2. Distribution of NSFW vs. SFW amongst the four offensive image detection models on the raw data. To ensure final labale assignment robustness, we utilize a voting mechanism to balance individual model biases.
We utilize the frame processed above and static images into the four models, where we obtain the nsfw to sfw ratio. We present the results in Figure 2. nsfw-image-detection is 0.88%, nsfw-classifier is 0.93%, vit-base-violence-detection is 27.27%, vit-base-nsfw-detector is 0.47%.
@@ 3.4.2. Unsafe Image Removal and Replacement
For the identified offensive images, we replace the textual content with the placeholder U-Sticker detects this as offensive image. This helps maintain the integrity of the dataset while ensuring that no harmful visuals, such as explicit or violent content, are retained, thus reducing the risk of privacy violations or inappropriate representations. Then, we remove the sticker entirely from our sticker database.
@@ 3.5.1. User Identifier
To protect the user’s privacy, we anonymize the user identifiers by hashing their original integer identifiers using the SHA-256 hash function (Appel, 2015). Originally, the user identifiers were integers ranging from five-digital value to a maximum value of ten-digit value. After applying the SHA-256 hashing algorithm (Appel, 2015), the user identifiers are transformed into 256-bit strings, which are irreversible and do not retain any direct correlation with the original numeric values. This ensures that the users’ identities are kept private, while still allowing for secure and effective data processing and analysis.